### v3

@@ 2 Literature review
Finally, this paper adds to the literature on the gender gap in the tech sector. Various reasons for low women’s participation in the technology sector have been studied. For example, Murciano-Goroff (2022) identify women’s tendency to under-report programming skills, impacting their job prospects. Strategies proposed to counter this include increasing role model visibility (Correll and Mackenzie (2016)) and shifting the perception of tech roles being predominantly male (Cheryan et al. (2013), Del Carpio and Guadalupe (2022)). We contribute to this literature by demonstrating how mentoring and job-training programs can enhance women’s success rates in obtaining tech jobs. Notably, our work focuses on the primary outcome of finding a job.
3 Mentoring and Challenges Programs
Over recent decades, Poland’s technology sector has witnessed significant growth, culminating in a 2021 revenue of roughly USD 51 billion, which accounts for 8% of the national GDP and provides employment for 250,000 individuals (GUS, 2022a). The sector’s growth has contributed to an escalating demand for skilled tech professionals, subsequently driving an approximate 9% wage growth in 2021. This led to the sector’s median monthly wage reaching USD 2,600, substantially outpacing the country’s overall median monthly wage for full-time workers, which stands at approximately USD 950 (GUS, 2022b). A survey of gender demographics within this industry unveils an underrepresentation of women, who make up 30% of the total Polish tech workforce (Gawlowska-Bujok and Bujok, 2021). The European Commission (European Commission, 2021) reports that the tech industry accounts for 1.1% of Polish women active in the labor market, compared to 5.2% of men.
Dare IT is a social impact organization that aims at reducing the gender gap in the tech sector. They do so by helping qualified women transition from other careers. Dare IT offers a variety of services for women looking for jobs in tech or making their first steps in the profession, including a mentoring program, career fairs, networking events, and an online social network.
@@ 3 Mentoring and Challenges Programs
Dare IT launched Mentoring in 2018 as a not-for-profit initiative to support women aspiring to tech jobs. Since then, the program has been offered twice per year, and the number of participants has increased in each cohort. The program is funded by grants and donations, which allows participation to be free of charge. In recent years, demand for the program has been far greater than supply. For example, for the spring 2021 cohort, Dare IT received over 1,000 applications for 150 spots.
The program follows several steps. Each cycle begins with Dare IT recruiting potential mentors through past participant referrals, social media engagement, and corporate partnerships. All mentors are women who are active professionals in the tech sector, and they contribute their time and expertise to the program on a pro-bono basis. Mentors are screened and selected to participate based on an application process; Dare IT considers multiple criteria, including candidates’ expertise in domains that attract the most interest from mentees, length of professional experience, and reputation.
Simultaneously, the recruitment of potential mentees occurs primarily via social media and word-of-mouth recommendations. During this study, Dare IT did not utilize paid advertisements for recruitment. Dare IT selects participants based on several criteria. Only fully-completed applications by women not currently engaged in a tech role are considered. The program primarily targets individuals with necessary skills — quantified by self-reported learning hours in a particular domain and manual review of applications by Dare IT employees — who require additional support in their tech job search. Applicants who do not meet a baseline skill level are excluded. After Dare IT finalizes the selection process, mentors receive access to a database of eligible applicants. They can then select their preferred mentee.
Once mentors and mentees are paired, they typically meet on a weekly basis for four months in order to support the mentee in their job search. Dare IT provides mentors with guidelines and best practices to increase the likelihood of a successful mentorship experience, although mentors have the flexibility to modify these guidelines to cater to their mentee’s needs. As a result, mentoring takes various forms, from coaching in job searching and interviewing or resume preparation to reviewing code and developing technical skills. Our discussions with mentors and Dare IT staff suggest four potential channels through which Mentoring increases the employability of mentees: (i.) reinforcement of self-efficacy through the provision of emotional support and encouragement, (ii.) help in the development of a professional network by relaying on connections of mentors [^7] Pallais and Sands (2016) emphasizes how referrals can augment job opportunities and delves into the underlying mechanisms., (iii.) explanation of unwritten norms of professional behavior and conduct, especially those intrinsic to the tech sector and, more specifically, during tech recruitment phases, and (iv.) direct transfer of skills through code review and discussions.
Challenges.
Due to the excess demand for Mentoring, Dare IT explored the possibility of expanding its capacity. However, scaling proved to be a complex issue. First, selecting more mentors without sacrificing the quality of mentoring is difficult due to the scarcity of women occupying suitable roles in tech. Second, implementing a large 1:1 mentoring program presents operational difficulties.
We partnered with Dare IT to develop a more easily scalable approach. Our first step was forming a set of hypotheses on barriers that impede women from securing roles in the Polish tech sector. We conducted interviews with hiring managers and HR experts from a dozen prominent tech companies in Poland. Each of our interviewees mentioned a shortage of practical experience or signals of practical skills on job application materials provided by female candidates. They emphasized that tangible proof of hands-on experience, even a short project, could significantly alter the interview dynamics, typically favoring candidates. They also expressed low trust in skill credentials that are available online. Many of our interviewees described online credentials as needing to be more informative about the capability to carry out real-world business tasks. Motivated by these findings, we developed a new program called Challenges, designed to address these specific concerns.
The Challenges program comprises six bi-weekly mini-projects. Successful completion of all mini-projects results in a portfolio item, e.g., a UX design for a mobile application, which functions as proof of practical experience in the chosen domain (see Figure 1 for an example of wireframes from a completed product). The program strongly emphasizes on-the-job training: assignments are curated by domain experts and are readily applicable in a business setting. Each weekly mini-project undergoes evaluation and feedback from domain practitioners, including existing mentors and industry professionals, facilitating ongoing skill and knowledge enhancement for participants. A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
Figure 1: Challenges final product.
We launched the program in December 2021, setting the start date for the initial cohort in mid-January 2022. Dare IT promoted the program on its social media channels to encourage applications. Enrollment was capped at 300 participants. [^8] The decision to limit program capacity to 300 spots was made to account for potential logistical challenges and to ensure that we could provide personalized assistance to all participants if necessary. As the program progressed, it became evident that only a few participants required hands-on support. Consequently, in subsequent cohorts, Dare IT increased the number of available spots. The first 100 slots were allocated to the initial 100 applicants (excluded from this analysis), with the remaining 200 spots randomly assigned among subsequent applicants. The program was oversubscribed within four hours, at which point we closed the registration form. Applications received post-closure were not included in the randomization. However, we collected data on these late applicants’ outcomes for comparison.
4 Experimental design and data
4.1 Experimental design
Mentoring.
We evaluate the mentoring program using the cohort of Autumn 2021. As in previous cohorts, Dare IT received a significantly higher number of applications than the available number of mentors. Even after the screening process to verify eligibility, the ratio of applicants to mentors remained at 4:1. The eligibility for the Mentoring program underwent manual verification by Dare IT staff. Only applicants who did not currently hold positions in the technology field and who, according to Dare IT’s reviewer assessments, have technical skills at a level typical of a junior employee in a tech company were considered eligible for participation. The surplus of eligible applicants compared to the number of available mentors provided us with an opportunity to implement a randomized access experiment. In order to maintain high-quality mentor-mentee pairings while incorporating randomized access, we asked mentors to select two potential mentees instead of just one, as they did in the previous program’s cohorts. Within these pairs, we then randomly determined which of the two applicants would be granted access to the program (the treatment group) and which would not (the control group). Thus, we created a paired experiment.
In the experiment, we initially had 156 mentor-mentee pairs and an equivalent number of mentor-control pairs. However, four pairs did not start the program because the mentee secured a tech job between applying for the program and its commencement. In these cases, the mentor-selected applicant, who was initially randomized into the control group, participated in the program. [^9] We decided against having idle mentors to ensure broader access to the program, as well as, to respect mentors’ time dedicated to selecting candidates and willingness to volunteer their time in mentoring. Consequently, we excluded these mentors and participants from our analysis. As a result, our analysis included a total of 152 pairs. Additionally, we gathered data on 468 eligible applicants whom mentors did not select for participation.
Challenges.
Regarding the Challenges program, we concluded the recruitment process once we received 500 applications that met the established eligibility criteria. To qualify for participation, applicants were required to declare that they were not currently employed in the technology sector, possess a fluency level in English (B1 level), and be proficient in Figma or React depending on the specialization. The first 100 applicants were granted program spots, while the remaining 400 were randomly assigned to either the treatment or control groups. Randomization was conducted with stratification based on variables obtained from the registration survey. [^10] These stratification variables included specialization (front-end development or UX), age, type of residence, and employment status (student, employed, neither employed nor student). Additionally, we gathered data on 160 applicants who submitted their forms after the registration process had concluded. These late applicants were not included in the primary program evaluation. However, they are considered in the off-policy evaluation, as detailed in Section 6.
Furthermore, it’s worth noting that some applicants randomized into the treatment group did not actively participate in the program (did not submit any assignments). Nevertheless, since these individuals still received all program materials and had the opportunity to engage in webinars discussing projects, we have included all such individuals in our analysis.
@@ 4.2 Data
We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted following the program’s conclusion, and information collected from participants’ LinkedIn profiles.
@@
Applicants in both programs were asked to complete a simple registration survey. Data from the survey included the city of residence, age, employment situation, and time spent, which measures how much time the applicant had spent developing their IT skills. The mentoring application form was richer; first, applicants were asked if they have family or friends in IT and whether they have children; second, the form also had several open-ended questions related to the applicant’s motivation and the possibility of attaching a short introductory video, but we do not use either in this analysis. Additionally, we collected information on participants’ preferred domains: the main categories were UX design, manager (project or product), or developer (front or back-end).
@@
Mentors were asked a number of questions before the start of the program, which allowed us to construct the following variables: first-time mentor, career changer (which indicates whether the mentor herself had worked outside of IT before moving to the sector), managerial experience (which indicates whether the mentor has any managerial experience or not), and years of experience (years of experience in the tech sector).
Outcomes survey.
We surveyed participants of each program and those in control groups four months after the end of the program. In the survey, we were primarily interested in the salaries in the new jobs: we asked the question whether the respondent found a new job and in which range the salary was. We also inquired about the number of job offers and whether the participant negotiated the proposed offer or accepted the initial one. Finally, we carried out a shorter survey right after the end of Mentoring, the “short survey,” where we asked for salaries for the new jobs. Altogether we conducted three surveys: the main mentoring survey in which we received 162 responses (with 111 responded from treatment group subjects), the initial survey of the mentoring program (with 194 responses), and the Challenges survey, for which we received responses from 68 subjects in the control group and 61 from the treatment group. We analyze the outcomes from surveys in Appendix F.
LinkedIn data.
Our primary source of outcome data is the LinkedIn profiles of the participants. This data was collected at two different points: first, four months, and second, 12 months after the conclusion of the programs. During the initial data collection phase, outcomes for all experimental subjects were gathered. In the second phase, Dare IT was unable to observe the outcome of one Mentoring participant, 17 Challenges participants, and 7 Challenges control group members. The principal cause of the missing data is the change in the LinkedIn profile address. [^11] It’s plausible that the propensity to change LinkedIn’s profile address was impacted by participation in the program. In Appendix ate_bounds we provide robustness checks to our primary results by assuming the most pessimistic and optimistic outcomes for the missing subjects. We also estimate AT bounds following the methodology in Lee and Weidner (2021). The primary outcome metric was whether a participant secured a technology job following a program’s start. As new technology jobs, we consider jobs with a starting date of January 2022 or later for Mentoring and February 2022 or later for Challenges that fall into one of the following two categories:
@@
    All jobs in technology companies other than positions in finance, regulatory, legal, accounting, and HR, where technology companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online shops).
@@
    Jobs in non-technology companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
Based on these categories, we define three outcomes: new job, tech job 4 months, and tech job 12 months. New job refers to any job, whether in technology or not, that started between January and May 2022 for Mentoring and between February and June for Challenges. Tech job 4 months denotes a job that aligns with category (1) or (2) and is also classified as a new job. Finally, tech job 12 months is a job, corresponding to either category (1) or (2), that began after January 2022 for Mentoring and February for Challenges and before May 2023. We also collect the date on which subjects got their new jobs. In some cases, we do not observe the month in which the new job started, but only the year; for such observations, we assign the start date as December 2022 or May 2023, depending on whether the job started in 2022 or 2023.
We also leveraged LinkedIn to gather data on the highest education level attained by participants, which was organized into the following categories: bachelor’s degree, master’s degree, postgraduate degree, non-traditional credential, and other. The other category includes profiles lacking education information or information that could not be classified into any of the pre-existing categories. Furthermore, we compiled data on college majors, forming two binary variables: STEM and social science or business. The former includes all STEM degrees, excluding social sciences, while the latter comprises degrees in social science, business, finance, or management. Lastly, we deduced a person’s age from a combination of age reported in the registration survey and the commencement date of their latest educational degree. Finally, we also compute their professional experience length by subtracting the start date of their first job recorded on LinkedIn from 2022. Table 1 summarizes the key variables for both programs.
@@
First time mentor 152 0.467 0.501 0 1 145 0.434 0.497 0 1
Managerial experience 152 0.520 0.501 0 1 145 0.510 0.502 0 1
Career changer 152 0.691 0.464 0 1 145 0.710 0.455 0. 1
Years of tech experience 152 5.856 3.777 2 35 145 6.021 3.794 2 35
@@
New job 152 0.467 0.501 0 1 147 0.422 0.496 0 1
Tech job 4 months 152 0.421 0.495 0 1 147 0.293 0.456 0 1
Tech job 12 months 151 0.609 0.499 0 1 147 0.456 0.500 0 1
Social science 152 0.132 0.339 0 1 147 0.136 0.344 0 1
STEM 152 0.487 0.501 0 1 147 0.531 0.501 0 1
Family friends IT 152 0.697 0.461 0 1 147 0.701 0.460 0 1
Warsaw 152 0.276 0.449 0 1 147 0.259 0.439 0 1
Mother 152 0.257 0.438 0 1 147 0.259 0.439 0 1
Applied before 152 0.243 0.431 0 1 147 0.231 0.423 0 1
Years of professional experience 152 7.053 5.049 0 31 147 7.204 4.816 0 23
@@
New job 183 0.333 0.473 0 1 225 0.293 0.456 0 1
Tech job 4 months 183 0.284 0.452 0 1 225 0.196 0.398 0 1
Tech job 12 months 166 0.380 0.487 0 1 218 0.266 0.443 0 1
Social science 183 0.208 0.407 0 1 225 0.267 0.443 0 1
STEM 183 0.421 0.495 0 1 225 0.391 0.489 0 1
Years of professional experience 183 7.645 4.976 0 31 225 7.671 6.237 0 31
@@
Tech job 4 months - - - - - 628 0.152 0.359 0 1
Tech job 12 months - - - - - 628 0.279 0.449 0 1
STEM - - - - - 628 0.196 0.367 0 1
Years of professional experience - - - - - 628 7.474 6.318 0 37

### v4

@@ 2 Literature Review
Finally, this paper adds to the literature on the gender gap in the technology sector. The persistent gender wage gap reflects both occupational segregation and within-occupation differences (juhn2017specialization). Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (murciano2022missing) and the impact of occupational stereotypes (cheryan2013stereotypical; del2022more). Strategies proposed to counter these barriers include increasing role model visibility (correll2016succeed) and shifting perceptions of technology roles. We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.
3 Experimental Interventions and Theoretical Framework
Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities. Workers address these through skill signaling (spence1978job) and professional networks (granovetter1973strength; ioannides2004job; pallais2016referential). For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers. First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (cheryan2013stereotypical; del2022more; murciano2022missing). Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (greguletz2019women; forret2004networking). Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness. Women systematically under-report their technical abilities on resumes compared to similarly skilled men (murciano2022missing), use less self-promoting language in professional contexts (lerchenmueller2019gender), and may experience reduced confidence when confronting occupational stereotypes (spencer1999stereotype). We study the relative importance of these barriers by evaluating two interventions with distinct mechanisms: a mentoring program that addresses network gaps and builds confidence through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
Context.
Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (gus2022). Despite strong labor demand driving 9% annual wage growth, women held only 30% of technology positions and comprised just 1.1% of the female labor force, compared to 5.2% of men (EC2021; fluff2021). This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (unitednations2021).
3.1 Program Designs and Mechanisms
We partnered with Dare IT, a Polish social impact organization addressing gender barriers in technology employment, to evaluate two interventions targeting different constraints. Both programs serve women with baseline technical skills who have not yet secured technology jobs. The programs’ contrasting designs allow us to contrast network effects from signaling effects.
Mentoring: Network Expansion and Tacit Knowledge.
Dare IT launched its Mentoring program in 2018 to support women transitioning into technology careers. The program pairs participants with experienced female technology professionals for weekly one-on-one sessions over four months. Mentors, all women actively working in tech, volunteer their time and are selected based on technical expertise and professional standing. From its inception, demand has substantially exceeded capacity. The spring 2021 cohort received over 1,000 applications for 150 available spots, a ratio typical of recent cohorts. This persistent oversubscription reflects a fundamental scalability constraint: the program’s reliance on volunteer mentors limits expansion, as the pool of qualified female tech professionals willing and able to dedicate four months of weekly mentoring remains finite.
Mentoring addresses gender-specific network barriers documented in prior research. Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (greguletz2019women). Female professionals may also hesitate to engage in networking behaviors that appear effective for men (forret2004networking). By connecting mentees to established female professionals, the program potentially overcomes both barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
Our qualitative interviews with mentors reveal four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review. Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
Challenges: Skill Signaling.
To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes. Program development began with structured interviews of hiring managers at twelve prominent Polish technology firms. These conversations revealed a systematic pattern: hiring managers perceived female applicants as lacking practical technical experience, even when formal credentials suggested otherwise. Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications. These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (murciano2022missing) and stereotypes about women’s technical competence create additional screening burdens for female candidates (cheryan2013stereotypical). A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
Challenges provides six bi-weekly assignments in User Experience Design (UX) or front-end programing in React. All assignments are designed by industry practitioners, that culminate in a complete portfolio project (UX design or React application; see Figure 1 for an example from the UX path). Each assignment replicates realistic business tasks and receives expert evaluation and feedback. Participants can add completed portfolios to resumes and reference them in interviews, creating a signal that hiring managers indicated they would find credible. The program operates online, accommodates cohorts of 300+ participants, and costs approximately \$15 per person.
Figure 1: An Example of Challenges Final Product.
A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
3.2 Theoretical Predictions and Empirical Strategy.
The programs’ differential mechanisms generate testable predictions about the treatment effect heterogeneity. The patterns of heterogeneity reveal which populations benefit most from network expansion versus skill signaling, providing evidence on the economic importance of each mechanism.
If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts. The network mechanism also predicts effects on observable network measures: in particular, we should observe increases in LinkedIn connection counts among Mentoring participants. Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics. Mentors with longer technology sector tenure and those in managerial positions likely possess richer professional networks and better access to hiring decision-makers, potentially generating larger effects for their mentees.
If imperfect information about technical abilities constrains hiring, particularly when exacerbated by gender stereotypes, Challenges should disproportionately benefit women whose traditional credentials provide weak or ambiguous signals: for example, those early in their careers with limited work history. For these candidates, the portfolio complements weak existing signals. In contrast, women with strong observable credentials may gain less marginal value from an additional portfolio signal.
We test these predictions by estimating treatment effects across participant subgroups defined by network strength, credential quality, and demographic characteristics. Differential patterns of heterogeneity provide evidence on whether these frictions meaningfully constrain employment and which populations each intervention serves most effectively. We further examine effects on LinkedIn connection counts to directly test the network expansion mechanism.
4 Experimental Design and Data
@@
We evaluate the Autumn 2021 cohort, for which recruitment opened in October 2021 and the program launched in January 2022. As in previous cohorts, applications substantially exceeded the number of available mentors. Even after Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking technical skills at a junior professional level, the applicant-to-mentor ratio remained 4:1.
This oversubscription allowed us to implement a randomized evaluation while preserving Dare IT’s mentor-mentee matching process. We modified the standard selection procedure: instead of each mentor selecting one mentee (the historical practice), we asked mentors to select their two preferred candidates from the pool of eligible applicants. We then randomized within each pair, assigning one applicant to treatment (participation in Mentoring) and one to control (no participation). This paired randomization design controls for mentor selection preferences while generating exogenous variation in program access.
Of 160 mentors, 151 agreed to this modified selection process; 9 declined, because they did not want to spend effort selecting mentee who will not end up working with them. This yielded an initial sample of 302 applicants (151 treatment, 151 control) across 151 mentor pairs. However, 4 treatment-assigned participants secured technology jobs between randomization (November 2021) and program launch (January 2022). We excluded these individuals from analysis and allowed the initially-assigned control applicant to participate instead, maintaining program access for mentors who invested time in candidate selection. Our final experimental sample comprises 151 treatment participants and 147 control participants.
Appendix G provides robustness checks: (i.) excluding the 4 mentor pairs without matched controls, (ii.) including the 4 pre-program employment cases with outcomes set to 1, and (iii.) including the 9 non-participating mentors under the assumption their hypothetical mentees did not find technology jobs. Results are substantively unchanged across specifications.
In addition to experimental subjects, we collected outcome data on 468 eligible applicants whom mentors did not select. These applicants provide a benchmark for assessing mentor selection effects independent of program participation. As shown in Figure 2, selected-but-control participants have substantially higher employment success than non-selected applicants, highlighting the importance of randomization for isolating treatment effects from selection effects.
Challenges. We launched recruitment in December 2021, with the program scheduled to begin in February 2022. To generate initial interest, we guaranteed admission to the first 100 applicants. Registration was capped at 500 applicants, but high demand led to 525 applications before we closed the portal.
Applicants self-certified that they (i.) were not currently employed in the technology sector, (ii.) possessed English language proficiency at B1 level or higher, and (iii.) had working knowledge of either Figma (User Experience Design track) or React (Front-end Development track). Unlike Mentoring, we did not manually verify technical skills prior to admission due to scalability considerations.
We exclude the first 100 guaranteed-admission applicants from our analysis, as they were not subject to randomization. Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access). Randomization employed stratification on four registration characteristics: track specialization (UX Design vs. Front-end Development), age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
Between randomization (December 2021) and program launch (February 2022), 17 treatment-assigned applicants secured technology employment. Consistent with our approach in Mentoring, we exclude these individuals from our primary analysis. Our final analytic sample includes 183 treatment participants and 225 control participants. Appendix G demonstrates that results are robust to including these 17 early-employment cases with outcomes coded as 1.
Among the 183 treatment participants, 41 (22%) did not submit any program assignments, suggesting passive non-compliance. However, these individuals retained access to all program materials, weekly webinars with industry experts, and the Dare IT Slack community. Following intention-to-treat principles, we include them in our analysis regardless of active participation level.
We also collected outcome data on 160 applicants who submitted forms after enrollment closed. These late applicants were not randomized and do not contribute to our primary treatment effect estimates. However, they provide an additional comparison group for the off-policy evaluation in Section 6.
Appendix B reports covariate balance tests across treatment and control groups. Of 9 baseline characteristics, only social science degree shows a statistically significant imbalance at the 5% level, consistent with chance given multiple comparisons. Our main specifications control for all observed variables to improve precision.
@@ 4.1 Data
We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted four months post-program, and LinkedIn profile data.
@@
Both programs collected basic information during registration: city of residence, age, employment situation, and time spent developing IT skills. The Mentoring application additionally asked whether applicants have family or friends in IT and whether they have children. We also collected information on preferred domains: UX design, manager (project or product), or developer (front or back-end).
@@
Mentors provided information enabling us to construct: first-time mentor (binary), career changer (whether the mentor worked outside tech before transitioning to the sector), managerial experience (binary), and years of experience in the tech sector.
LinkedIn Data. Our primary outcome measure, pre-registered prior to data collection, is whether participants secured technology-sector employment following program completion. We collected employment data from participants’ LinkedIn profiles at two time points: four months post-program (May 2022 for Mentoring, June 2022 for Challenges) and twelve months post-program (May 2023 and June 2023, respectively).
LinkedIn provides standardized job title metadata for each employment entry. The research team and research assistants, blinded to treatment assignment, coded each profile according to our pre-registered definition. A technology job is:
@@
    Any position in a technology company—firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms—excluding finance, regulatory, legal, accounting, and HR roles.
@@
    Technical positions in non-technology companies involving software development and testing, IT support, or data analytics (e.g., software engineers at banks, data analysts at consulting firms).
Self-employment, freelancing, and volunteer positions were not coded as new employment. Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference (p=0.9), supporting their equivalence.
Data collection and attrition.
We obtained complete LinkedIn data for all Mentoring participants at both measurement points. For Challenges, we collected data on all participants at four months and 384 of 408 participants (94%) at twelve months. Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves. Appendix G provides bounds and estimates under extreme attrition assumptions; main results remain statistically significant under all scenarios.
Outcome variables.
We define three binary measures: new job (any employment starting January 2022 or later for Mentoring, February 2022 or later for Challenges), tech job (technology-sector position per above criteria), and non-tech job (new employment outside technology). In a few cases, subjects reported only the start year of new employment on their LinkedIn profiles, we conservatively assign December 2022 or May 2023 as the start date.
We also collected from LinkedIn: highest education level (bachelor’s, master’s, postgraduate, non-traditional credential, or other), college major (constructing binary variables for STEM and social science/business), age (combining registration survey data with degree completion dates), and professional experience length (from first job start date to 2022). Table 1 summarizes key variables.
@@
First time mentor 151 0.47 0.50 0 1 145 0.43 0.50 0 1
Managerial experience 151 0.52 0.50 0 1 145 0.51 0.50 0 1
Career changer 151 5.86 3.79 2 35 145 6.02 3.79 2 35
Years of tech experience 151 0.70 0.46 0 1 145 0.71 0.46 0 1
@@
New job 151 0.65 0.48 0 1 147 0.57 0.50 0 1
Non-tech job 151 0.04 0.20 0 1 147 0.12 0.32 0 1
Tech job 151 0.61 0.49 0 1 147 0.46 0.50 0 1
LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
STEM 151 0.49 0.50 0 1 147 0.53 0.50 0 1
Student 151 0.14 0.35 0 1 147 0.16 0.37 0 1
Over 30 151 0.47 0.50 0 1 147 0.50 0.50 0 1
Small town 151 0.25 0.44 0 1 147 0.24 0.43 0 1
Applied before 151 0.24 0.43 0 1 147 0.23 0.42 0 1
Mother 151 0.26 0.44 0 1 147 0.26 0.44 0 1
High proficiency 151 0.54 0.50 0 1 147 0.57 0.50 0 1
Family or friends in IT 151 0.70 0.46 0 1 147 0.70 0.46 0 1
UX 151 0.42 0.50 0 1 147 0.44 0.50 0 1
Masters 151 0.48 0.50 0 1 147 0.41 0.49 0 1
Years of professional experience 151 6.99 5.01 0 31 147 7.20 4.82 0 23
@@
New job 166 0.57 0.50 0 1 218 0.50 0.50 0 1
Non tech job 166 0.19 0.40 0 1 218 0.23 0.42 0 1
Tech job 166 0.38 0.49 0 1 218 0.27 0.44 0 1
LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
STEM 183 0.42 0.50 0 1 225 0.39 0.49 0 1
Social science 183 0.21 0.41 0 1 225 0.27 0.44 0 1
Warsaw 183 0.20 0.40 0 1 225 0.20 0.40 0 1
Years of professional experience 183 7.64 4.98 0 31 225 7.67 6.24 0 49
UX 183 0.49 0.50 0 1 225 0.54 0.50 0 1
Over 30 183 0.47 0.50 0 1 225 0.44 0.50 0 1
Small town 183 0.61 0.49 0 1 225 0.63 0.48 0 1
@@
Tech job 12 months - - - - - 628 0.28 0.45 0 1
STEM - - - - - 628 0.20 0.40 0 1
Warsaw - - - - - 628 0.31 0.46 0 1
Small town - - - - - 628 0.39 0.49 0 1
Years of professional experience - - - - - 606 7.47 6.32 0 37
Over 30 - - - - - 628 0.14 0.35 0 1
Masters degree - - - - - 628 0.31 0.46 0 1
Mentoring applicant - - - - - 628 0.75 0.44 0 1
@@
Outcomes Survey.
We surveyed participants and control group members four months post-program, asking about new employment, salary ranges, number of job offers, and negotiation behavior. We conducted three surveys: the main Mentoring survey (162 responses, 111 from treatment), an initial Mentoring survey (194 responses), and the Challenges survey (68 control, 61 treatment). Appendix E analyzes survey outcomes.